OpenStudy (anonymous):

I need help,Please! An pathropologist studying the bones of a prehistoric person finds there is so little remaining Carbon-14 in the bones that instruments cannot mesure it .This mean that there is less than 0.5% of the amount of the amount of Carbon-14 the bones would have contained wnen the person was alive .How long ago did person die? the key is more than 44.000 years ago Page 548

OpenStudy (anonymous):

The key is not "more than 44.000 years ago." The key is to find, through a straightforward google search, the half-life of C-14, which is 5730 years. We have A(t)=A0*(1/2)^(t/5730) 0.0005A0=A0*(1/2)^(t/5730) (1/2)^(t/5730) = 0.0005 Take the ln of both sides: (t/5730)ln(1/2) = ln(0.0005) t/5730 = 10.9658 t is approximately 62,833.94 (years ago).

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

can you help me one more?

OpenStudy (anonymous):

I will post another page